OpenStudy (anonymous):

Find the measure of angle x. Round your answer to the nearest hundredth. (please type the numerical answer only) A right triangle is shown with one leg measuring 12 and the hypotenuse is labeled 13. An angle that is not 90 degrees is marked x.

OpenStudy (anonymous):

adjacent is 12

OpenStudy (aum):

There are two angles that are not 90 degrees. Which angle is marked "x" in the diagram? Or you can simply post a screenshot of the diagram or draw one yourself using the "Draw" button below the reply box.

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@aaronq

OpenStudy (aum):

The angle "x" is formed by an adjacent side and a hypotenuse. Cosine is a trig function that relates the adjacent side to the hypotenuse. cos(x) = adjacent / hypotenuse = 12/13 Angle x = inverse cosine of (12/13) = ? Use a calculator to find the inverse cosine.